You will receive 32 annual payments of $37,500. The first payment will be received 7 years from today and the interest rate is 6.6 percent. What is the value of the payments today?
Answer #1

N = 32

Annual payments = 37500

Rate = 6.6%

Value of payments in year 6 = A* (1-1/(1+r)^n)/r

= 37500*((1-1/1.066^32)/0.066)

=494,686.99

Present value of the payments = FV/ (1+r)^n

= 494,686.99/ 1.066^6

=337,122.13